# Ordered Momentum for Asynchronous SGD

### v2

@@ 1 Introduction
Many machine learning problems, such as support vector machine (SVM), logistic regression, and deep learning models, can be formulated as the following empirical risk minimization problem:
$\min_{{\bf w}\in{\mathbb{R}}^{d}}F({\bf w})=\frac{1}{|{\mathcal{D}}|}\sum_{\xi\in{\mathcal{D}}}f({\bf w};\xi),$ (1)
where ${\bf w}$ denotes the model parameter, $d$ is the dimension of the parameter, ${\mathcal{D}}$ represents the training dataset, $|{\mathcal{D}}|$ is the number of the training instances and $f({\bf w};\xi)$ denotes the empirical loss on the training instance $\xi$.
@@ 2 Preliminary
One of the most representative methods for distributing SGD across multiple workers is Synchronous SGD (SSGD) [20, 10]. Distributed SGD (DSGD), as presented in Algorithm 1, unifies SSGD and ASGD within a single framework [13]. The waiting set ${\mathcal{C}}$ in Algorithm 1 is a collection of workers (indexes) that are awaiting the server to send the latest parameter. The only difference between SSGD and ASGD is the communication scheduler associated with the waiting set. SSGD corresponds to DSGD with a synchronous communication scheduler, while ASGD corresponds to DSGD with an asynchronous communication scheduler. We use ${\bf g}_{ite(k_{t},t)}^{k_{t}}$ to denote the stochastic gradient $\nabla f({\bf w}_{ite(k_{t},t)};\xi^{k_{t}})$, where $k_{t}$ is the index of the worker whose gradient participates in the parameter update at iteration $t$. The function $ite(k,t)$ denotes the iteration index of the latest parameter sent to worker $k$ before iteration $t$, where $k\in[K]$ and $t\in[T]$. When $K=1$, DSGD degenerates to vanilla SGD, i.e., $ite(k_{t},t)\equiv t$. The delay of the gradient ${\bf g}_{ite(k_{t},t)}^{k_{t}}$ is defined as $\tau_{t}=t-ite(k_{t},t)$.
@@
2:  Input: number of workers $K$, number of iterations $T$, learning rate $\eta$, waiting set ${\mathcal{C}}=\emptyset$;
3:  Initialization: initial parameter ${\bf w}_{0}$;
@@
15:     Randomly sample $\xi^{k}$ and compute the stochastic gradient ${\bf g}^{k}=\nabla f({\bf w};\xi^{k})$;
@@ 3.1 Reformulation of SSGD with Momentum
where $\tilde{{\bf u}}_{0}={\bf 0},\beta\in[0,1),s\in[S]$ and $S$ denotes the number of iterations. $\tilde{{\bf u}}_{s}$ represents the Polyak’s momentum. $\frac{1}{|{\mathcal{B}}_{s}|}\sum_{\xi\in{\mathcal{B}}_{s}}\nabla f\left(\tilde{{\bf w}}_{s};\xi\right)$ denotes the stochastic gradient computed over the sampled training instance set ${\mathcal{B}}_{s}$, which contains either a single training instance or a mini-batch of training instances sampled from ${\mathcal{D}}$. (3) denotes the parameter update step and (4) denotes the momentum update step. When $\beta=0$, SGDm degenerates to (mini-batch) SGD.
@@ 3.2 OrMo for ASGD
2:  Input: number of workers $K$, number of iterations $T$, learning rate $\eta$, momentum coefficient $\beta\in[0,1)$, waiting set ${\mathcal{C}}=\emptyset$;
3:  Initialization: initial parameter ${\bf w}_{0}$, momentum ${\bf u}_{0}={\bf 0}$, index of the latest gradient group $I_{0}=0$;
@@
21:     Randomly sample $\xi^{k}$ and calculate the stochastic gradient ${\bf g}_{t^{\prime}}^{k}=\nabla f({\bf w}_{t^{\prime}};\xi^{k})$;
@@
On the contrary, the sequence of gradients computed in ASGD is given by:
@@
We use $I_{t+1}$ to denote the index of the latest gradient group of ${\bf u}_{t+1}$ in OrMo. The iteration index of the latest gradient in ${\bf u}_{t+1}$ can be $t$ at most. Since the gradient with iteration index $t$ belongs to the $\lceil\frac{t}{K}\rceil$-th gradient group, the latest gradient group for ${\bf u}_{t+1}$ should be the $\lceil\frac{t}{K}\rceil$-th gradient group, i.e., $I_{t+1}\equiv\lceil\frac{t}{K}\rceil,\forall t\in[T]$. ${\bf u}_{t+1}$ is the weighted sum of the gradients from the first $I_{t+1}+1$ gradient groups, where some gradients may be missing because they have not yet arrived at the server. The gradients in the $i$-th gradient group are weighted by $\beta^{\lceil\frac{t}{K}\rceil-i}$ in the momentum ${\bf u}_{t+1}$, where $i\in[I_{t+1}+1]$. Figure 2 shows an example of ${\bf u}_{10}$ in OrMo when $K=4$.
@@
    The update rule of the parameter in OrMo is motivated by that in SSGDm, as presented in Algorithm 3. In SSGDm, the scaled gradient $\eta{\bf g}_{ite(k_{t},t)}^{k_{t}}$ is always added to the momentum with a weight of $\beta^{0}$. At the current iteration, this scaled gradient updates the parameter with a coefficient $-1.0$. In subsequent iterations, this scaled gradient in the momentum updates the parameter with the coefficients $-\beta,-\beta^{2},-\beta^{3},\cdots$. By the time this scaled gradient is weighted by $\beta^{I_{t+1}-\lceil\frac{ite(k_{t},t)}{K}\rceil}$ in the momentum of SSGDm, it has already updated the parameter for $I_{t+1}-\lceil\frac{ite(k_{t},t)}{K}\rceil+1$ steps, with a total coefficient $-\sum_{j=0}^{I_{t+1}-\lceil\frac{ite(k_{t},t)}{K}\rceil}\beta^{j}$. In OrMo, for the scaled gradient $\eta{\bf g}_{ite(k_{t},t)}^{k_{t}}$ which is added to the momentum with a weight of $\beta^{I_{t+1}-\lceil\frac{ite(k_{t},t)}{K}\rceil}$, we compensate for the missed $I_{t+1}-\lceil\frac{ite(k_{t},t)}{K}\rceil+1$ steps compared with SSGDm and update the parameter with the coefficient $-\frac{1-\beta^{I_{t+1}-\lceil\frac{ite(k_{t},t)}{K}\rceil+1}}{1-\beta}$ at the current iteration. The design of the parameter update rule is crucial for the derivation of Lemma 2, which is further supported by the ablation study in Appendix A.3.
@@ 3.3 Convergence Analysis
In this section, we prove the convergence of OrMo in Algorithm 2 for non-convex problems. We only present the main results here. The proof details can be found in Appendix C.2.
@@
For any stochastic gradient $\nabla f({\bf w};\xi^{k})$, we assume that it satisfies:
${\mathbb{E}}_{\xi^{k}}[\nabla f({\bf w};\xi^{k})]=\nabla F({\bf w}),\quad{\mathbb{E}}_{\xi^{k}}\|\nabla f({\bf w};\xi^{k})-\nabla F({\bf w})\|^{2}\leq\sigma^{2},\quad\forall{\bf w}\in{\mathbb{R}}^{d},\forall k\in[K].$
@@
For any stochastic gradient $\nabla f({\bf w};\xi^{k})$, we assume that it satisfies:
$\mathbb{E}_{\xi^{k}}\|\nabla f({\bf w};\xi^{k})\|^{2}\leq G^{2},\forall{\bf w}\in{\mathbb{R}}^{d},\forall k\in[K].$
@@
Assumption 1 indicates that the data distributions across all workers are independent and identically distributed (i.i.d.). It’s commonly satisfied in the data-center scenario for distributed training [5], where all workers have access to the full dataset. Our analysis based on Assumption 1 also provides insights into the analysis in a non-i.i.d. setting [22], which will be studied in future work.
@@
Many works [46, 30, 7, 13, 22] consider delay-adaptive methods for ASGD. The key insight of these delay-adaptive methods is to penalize the gradients with large delays and reduce their contribution to the parameter update. OrMo is orthogonal to these delay-adaptive methods. Concretely, we can replace the constant learning rate $\eta$ in Algorithm 2 with a delay-adaptive learning rate $\eta_{t}$, which is dependent on the delay of the gradient $\tau_{t}$. Inspired by [13], we adopt the following delay-adaptive learning rate $\eta_{t}$:
@@
$\mathbb{E}\left\|\nabla F(\tilde{{\bf w}}_{T})\right\|^{2}\leq\mathcal{O}\left(\sqrt{\frac{L\sigma^{2}}{T}}+\frac{KL}{T}\right),$
where $\Delta=F({\bf w}_{0})-F^{*}$ and $\tilde{{\bf w}}_{T}$ is randomly chosen from $\{{\bf w}_{0},{\bf w}_{1},\cdots,{\bf w}_{T-1}\}$ according to (17).
The proof details can be found in Appendix C.3. Compared with Theorem 1, Theorem 2 removes the dependence on the bounded gradient assumption in Assumption 2 and provides a better convergence bound.
@@
Most existing theoretical analyses of ASGD [16, 49, 3, 20] rely on the maximum delay $\tau_{max}$ (e.g., ${\mathcal{O}}(\sqrt{\frac{L\sigma^{2}}{T}}+\frac{\tau_{max}L}{T})$ in [20]). However, since ASGD can still perform well even when the maximum delay is extremely large ($\tau_{max}\gg K$) in practice, these theoretical analyses don’t accurately reflect the true behavior of ASGD. The most closely related works to this work are [13, 22], which analyze ASGD without relying on the maximum delay. But the works in [13, 22] do not consider momentum. To the best of our knowledge, this is the first work to establish the convergence guarantee of ASGD with momentum without relying on the maximum delay.
@@ Appendix B Algorithm Details
21:     Randomly sample $\xi^{k}$ and calculate the stochastic gradient ${\bf g}^{k}=\nabla f({\bf w};\xi^{k})$;
@@
16:     Randomly sample $\xi^{k}$ and calculate the stochastic gradient ${\bf g}^{k}=\nabla f({\bf w};\xi^{k})$;
@@ C.2.3 Proof of Lemma 3
$\leq\frac{\beta^{2}q_{t}}{(1-\beta)^{2}}\left(\sum_{k\in[K]}\beta^{\lfloor\frac{t+K-2}{K}\rfloor}\left\|\eta{\bf g}_{0}^{k}\right\|^{2}+\sum_{s=1}^{\lfloor\frac{t+K-2}{K}\rfloor}\sum_{j=(s-1)K+1}^{\min\{sK,t-1\}}\beta^{\lfloor\frac{t+K-2}{K}\rfloor-s}\left\|\eta{\bf g}_{j}^{k_{j-1}}\right\|^{2}\right)$
@@ C.2.4 Proof of Theorem 1
$\frac{1}{T}\sum_{t=1}^{T}{\mathbb{E}}\|\nabla F({\bf w}_{t})\|^{2}$ $\leq\frac{4(1-\beta)\left[{\mathbb{E}}F(\hat{{\bf y}}_{1})-F^{*}\right]}{T\eta}+\frac{2L\eta\sigma^{2}}{1-\beta}+\frac{4\eta^{2}K^{2}G^{2}L^{2}}{(1-\beta)^{4}}$
@@
$\leq F({\bf w}_{0})+\frac{KL\sigma^{2}\eta^{2}}{2(1-\beta)^{2}}$
@@
$\overset{T\geq K}{\leq}\frac{4(1-\beta)\left[F({\bf w}_{0})-F^{*}\right]}{T\eta}+\frac{4L\eta\sigma^{2}}{1-\beta}+\frac{4\eta^{2}L^{2}K^{2}G^{2}}{(1-\beta)^{4}}$
@@ C.3.1 Algorithm
3:  Initialization: initial parameter ${\bf w}_{0}$, momentum ${\bf u}_{0}={\bf 0}$, index of the latest gradient group $I_{0}=0$;
@@
21:     Randomly sample $\xi^{k}$ and calculate the stochastic gradient ${\bf g}_{t^{\prime}}^{k}=\nabla f({\bf w}_{t^{\prime}};\xi^{k})$;

### v3

@@ 1 Introduction
Many machine learning problems can be formulated as optimization problems of the following form:
$\min_{{\bf w}\in{\mathbb{R}}^{d}}F({\bf w})={\mathbb{E}}_{\xi\sim{\mathcal{D}}}\left[f({\bf w};\xi)\right],$ (1)
where ${\bf w}$ denotes the model parameter, $d$ is the dimension of the parameter, ${\mathcal{D}}$ represents the distribution of the training instances and $f({\bf w};\xi)$ denotes the loss on the training instance $\xi$.
@@ 2 Preliminary
One of the most representative methods for distributing SGD across multiple workers is Synchronous SGD (SSGD) [20, 10]. Distributed SGD (DSGD), as presented in Algorithm 1, unifies SSGD and ASGD within a single framework [13]. The waiting set ${\mathcal{C}}$ in Algorithm 1 is a collection of workers (indexes) that are awaiting the server to send the latest parameter. The only difference between SSGD and ASGD is the communication scheduler associated with the waiting set. SSGD corresponds to DSGD with a synchronous communication scheduler, while ASGD corresponds to DSGD with an asynchronous communication scheduler. We use ${\bf g}_{ite(k_{t},t)}^{k_{t}}$ to denote the stochastic gradient $\nabla f({\bf w}_{ite(k_{t},t)};\xi^{k_{t}})$, where $k_{t}$ is the index of the worker whose gradient participates in the parameter update at iteration $t$ and $\xi^{k_{t}}$ denotes a training instance sampled on worker $k_{t}$. The function $ite(k,t)$ denotes the iteration index of the latest parameter sent to worker $k$ before iteration $t$, where $k\in[K]$ and $t\in[T]$. The delay of the gradient ${\bf g}_{ite(k_{t},t)}^{k_{t}}$ is defined as $\tau_{t}=t-ite(k_{t},t)$. When $K=1$, DSGD degenerates to vanilla SGD, i.e., $ite(k_{t},t)\equiv t$.
@@
2:  Input: number of workers $K$, number of iterations $T$, learning rate $\eta$;
3:  Initialization: initial parameter ${\bf w}_{0}$, waiting set ${\mathcal{C}}=\emptyset$;
@@
15:     Randomly sample $\xi^{k}\sim{\mathcal{D}}$ and then compute the stochastic gradient ${\bf g}^{k}=\nabla f({\bf w};\xi^{k})$;
@@ 3.1 Reformulation of SSGD with Momentum
where $\tilde{{\bf u}}_{0}={\bf 0},\beta\in[0,1),s\in[S]$ and $S$ denotes the number of iterations. $\beta$ is the momentum coefficient. $\tilde{{\bf u}}_{s}$ represents the Polyak’s momentum. $\frac{1}{|{\mathcal{B}}_{s}|}\sum_{\xi\in{\mathcal{B}}_{s}}\nabla f\left(\tilde{{\bf w}}_{s};\xi\right)$ denotes the stochastic gradient computed over the sampled training instance set ${\mathcal{B}}_{s}$, which contains either a single training instance or a mini-batch of training instances sampled from ${\mathcal{D}}$. (3) denotes the parameter update step and (4) denotes the momentum update step. When $\beta=0$, SGDm degenerates to (mini-batch) SGD.
@@ 3.2 OrMo for ASGD
3:  Initialization: initial parameter ${\bf w}_{0}$, momentum ${\bf u}_{0}={\bf 0}$, index of the latest gradient group $I_{0}=0$, waiting set ${\mathcal{C}}=\emptyset$;
@@
21:     Randomly sample $\xi^{k}\sim{\mathcal{D}}$ and then compute the stochastic gradient ${\bf g}_{t^{\prime}}^{k}=\nabla f({\bf w}_{t^{\prime}};\xi^{k})$;
@@
In contrast, the sequence of gradients computed in ASGD is given by:
@@
We use $I_{t+1}$ to denote the index of the latest gradient group of ${\bf u}_{t+1}$ in OrMo. The iteration index of the latest gradient in ${\bf u}_{t+1}$ can be $t$ at most. Since the gradient with iteration index $t$ belongs to the $\lceil\frac{t}{K}\rceil$-th gradient group, the latest gradient group for ${\bf u}_{t+1}$ should be the $\lceil\frac{t}{K}\rceil$-th gradient group, i.e., $I_{t+1}\equiv\lceil\frac{t}{K}\rceil,\forall t\in[T]$. ${\bf u}_{t+1}$ is the weighted sum of the gradients from the first $I_{t+1}+1$ gradient groups, where some gradients may be missing because they have not yet arrived at the server. The gradients in the $i$-th gradient group are weighted by $\beta^{\lceil\frac{t}{K}\rceil-i}$ in the momentum ${\bf u}_{t+1}$, where $i\in[I_{t+1}+1]$ and the superscript of the scalar $\beta$ indicates the exponent. Figure 2 shows an example of ${\bf u}_{10}$ in OrMo when $K=4$.
@@
    The update rule of the parameter in OrMo is motivated by that in SSGDm, as presented in Algorithm 3. In SSGDm, the scaled gradient $\eta{\bf g}_{ite(k_{t},t)}^{k_{t}}$ is always added to the momentum with a weight of $\beta^{0}$. At the current iteration, this scaled gradient updates the parameter with a coefficient $-\beta^{0}$. In subsequent iterations, this scaled gradient in the momentum updates the parameter with the coefficients $-\beta,-\beta^{2},-\beta^{3},\cdots$. By the time this scaled gradient is weighted by $\beta^{I_{t+1}-\lceil\frac{ite(k_{t},t)}{K}\rceil}$ in the momentum of SSGDm, it has already updated the parameter for $I_{t+1}-\lceil\frac{ite(k_{t},t)}{K}\rceil+1$ steps, with a total coefficient $-\sum_{j=0}^{I_{t+1}-\lceil\frac{ite(k_{t},t)}{K}\rceil}\beta^{j}$. In OrMo, for the scaled gradient $\eta{\bf g}_{ite(k_{t},t)}^{k_{t}}$ which is added to the momentum with a weight of $\beta^{I_{t+1}-\lceil\frac{ite(k_{t},t)}{K}\rceil}$, we compensate for the missed $I_{t+1}-\lceil\frac{ite(k_{t},t)}{K}\rceil+1$ steps compared with SSGDm and update the parameter with the coefficient $-\frac{1-\beta^{I_{t+1}-\lceil\frac{ite(k_{t},t)}{K}\rceil+1}}{1-\beta}$ at the current iteration. The design of the parameter update rule is crucial for the derivation of Lemma 2, which is further supported by the ablation study in Appendix A.3.
@@ 3.3 Convergence Analysis
In this section, we prove the convergence of OrMo in Algorithm 2 for non-convex problems. We only present the main results here. The proof details can be found in Appendix C.
@@
For any stochastic gradient $\nabla f({\bf w};\xi)$, we assume that it satisfies:
${\mathbb{E}}_{\xi\sim{\mathcal{D}}}\left[\nabla f({\bf w};\xi)\right]=\nabla F({\bf w}),{\mathbb{E}}_{\xi\sim{\mathcal{D}}}\|\nabla f({\bf w};\xi)-\nabla F({\bf w})\|^{2}\leq\sigma^{2},\forall{\bf w}\in{\mathbb{R}}^{d}.$
@@
For any stochastic gradient $\nabla f({\bf w};\xi)$, we assume that it satisfies:
$\mathbb{E}_{\xi\sim{\mathcal{D}}}\|\nabla f({\bf w};\xi)\|^{2}\leq G^{2},\forall{\bf w}\in{\mathbb{R}}^{d}.$
@@
Many works [46, 30, 7, 13, 22] consider delay-adaptive methods for ASGD. The key insight of these methods is to penalize the gradients with large delays and reduce their contribution to the parameter update. OrMo is orthogonal to these delay-adaptive methods. Concretely, we can replace the constant learning rate $\eta$ in Algorithm 2 with a delay-adaptive learning rate $\eta_{t}$, which is dependent on the delay of the gradient $\tau_{t}$. Inspired by [13], we adopt the following delay-adaptive learning rate $\eta_{t}$:
@@
$\mathbb{E}\left\|\nabla F(\bar{{\bf w}}_{T})\right\|^{2}\leq\mathcal{O}\left(\sqrt{\frac{L\sigma^{2}}{T}}+\frac{KL}{T}\right),$
where $\Delta=F({\bf w}_{0})-F^{*}$ and $\bar{{\bf w}}_{T}$ is randomly chosen from $\{{\bf w}_{0},{\bf w}_{1},\cdots,{\bf w}_{T-1}\}$ according to a probability distribution which is related to the delay-adaptive learning rates.
The proof details can be found in Appendix C.3. Compared with Theorem 1, Theorem 2 removes the dependence on Assumption 2 (bounded gradient) and provides a better convergence bound.
@@
We focus on the scenario where the training instances across all workers are independent and identically distributed (i.i.d.) from ${\mathcal{D}}$. This scenario commonly appears in the data-center setup for distributed training [5], where all workers have access to the full training dataset. Our analysis for the i.i.d. scenario can also provide insights into the analysis in a non-i.i.d. scenario [22], which will be studied in future work.
Remark 5.
Most existing theoretical analyses of ASGD [16, 49, 3, 20] rely on the maximum delay $\tau_{max}$ (e.g., ${\mathcal{O}}(\sqrt{\frac{L\sigma^{2}}{T}}+\frac{\tau_{max}L}{T})$ in [20]), where $\tau_{max}=\max_{t\in[T]}\tau_{t}$. However, since ASGD can still perform well even when the maximum delay is extremely large ($\tau_{max}\gg K$) in practice, these theoretical analyses don’t accurately reflect the true behavior of ASGD. The most closely related works to this work are [13, 22], which analyze ASGD without relying on the maximum delay. But the works in [13, 22] do not consider momentum. To the best of our knowledge, this is the first work to establish the convergence guarantee of ASGD with momentum without relying on the maximum delay.
@@ Appendix B Algorithm Details
21:     Randomly sample $\xi^{k}\sim{\mathcal{D}}$ and then compute the stochastic gradient ${\bf g}^{k}=\nabla f({\bf w};\xi^{k})$;
@@
16:     Randomly sample $\xi^{k}\sim{\mathcal{D}}$ and then compute the stochastic gradient ${\bf g}^{k}=\nabla f({\bf w};\xi^{k})$;
@@ C.2.3 Proof of Lemma 3
$\leq\frac{\beta^{2}q_{t}}{(1-\beta)^{2}}\left(\sum_{k\in[K]}\beta^{\lfloor\frac{t+K-2}{K}\rfloor}\left\|\eta{\bf g}_{0}^{k}\right\|^{2}+\sum_{s=1}^{\lfloor\frac{t+K-2}{K}\rfloor}\sum_{j=(s-1)K+1}^{\min\{sK,t-1\}}\beta^{\lfloor\frac{t+K-2}{K}\rfloor-s}\left\|\eta{\bf g}_{j}^{k_{j-1}}\right\|^{2}\right).$
@@ C.2.4 Proof of Theorem 1
$\frac{1}{T}\sum_{t=1}^{T}{\mathbb{E}}\|\nabla F({\bf w}_{t})\|^{2}$ $\leq\frac{4(1-\beta)\left[{\mathbb{E}}F(\hat{{\bf y}}_{1})-F^{*}\right]}{T\eta}+\frac{2L\eta\sigma^{2}}{1-\beta}+\frac{4\eta^{2}K^{2}G^{2}L^{2}}{(1-\beta)^{4}}.$
@@
$\leq F({\bf w}_{0})+\frac{KL\sigma^{2}\eta^{2}}{2(1-\beta)^{2}}.$
@@
$\overset{T\geq K}{\leq}\frac{4(1-\beta)\left[F({\bf w}_{0})-F^{*}\right]}{T\eta}+\frac{4L\eta\sigma^{2}}{1-\beta}+\frac{4\eta^{2}L^{2}K^{2}G^{2}}{(1-\beta)^{4}}.$
@@ C.3.1 Algorithm
3:  Initialization: initial parameter ${\bf w}_{0}$, momentum ${\bf u}_{0}={\bf 0}$, index of the latest gradient group $I_{0}=0$, waiting set ${\mathcal{C}}=\emptyset$;
@@
21:     Randomly sample $\xi^{k}\sim{\mathcal{D}}$ and then compute the stochastic gradient ${\bf g}_{t^{\prime}}^{k}=\nabla f({\bf w}_{t^{\prime}};\xi^{k})$;